# Trust, But Verify: A Survey of Randomized Smoothing Techniques

### v1

@@ III-A Randomized Smoothing for Training
The idea of incorporating randomized smoothing objectives in adversarial training [87] roots from the intuition that doing so will result in boosted performance and improved robustness. Further, addressing the challenge of training a good base classifier that is accurate and robust when smoothed [34] led to significantly better results as discussed in this section. Salman et al. [87] consider a smoothed classifier which is a generalization of equation (2), to soft classifiers, namely, functions $f:\mathbb{R}^{d}\rightarrow P(y)$ where $P(y)$ is the set of probability distributions over $y$. They then train a hard classifier which essentially takes the argmax of the soft classifier, using adversarial examples generated by describing an adversarial attack against the smoothed soft classifier [108, 109]. This results in a boosted empirical robustness and substantially improved certifiable robustness using the certification method of Cohen et al. [1]. Feng et al. [34] improved upon this by deriving a new regularized risk in which the regularization can ’adaptively encourage’ the training of the base classifier. Specifically, they proposed ADRE (ADaptive Radius Enhancing regularizer), which penalizes examples misclassified by the smoothed classifier while encouraging the certified radius of the correctly classified examples. It is computationally efficient and can be implemented in parallel with other empirical defense methods, under standard (non-adversarial) and adversarial training schemes. They also designed a new certification algorithm, T-CERTIFY, that can leverage the regularization effect to provide tighter robustness and lower bounds that hold high probability. Alfarra, Motasem, et al. [13] challenged these attack-dependent and time costly adversarial training techniques [1, 87] by introducing an algorithm called MACER, which learns robust models without using adversarial training. More specifically, they break the standard training loss into two subsequent losses: a standard cross-entropy loss and a hinge-based loss function applied to the robust radius. A modified version of the original randomized smoothing obtains the robust radius. Since it is a non-differentiable method, to make it differentiable, they take expectation over the probability scores each time a noisy sample is passed through the model and then take the maximum over all the classes. It is observed that MACER also outperforms Feng, Huijie, et al. [34] by a significant margin in terms of Certified top-1 accuracy on CIFAR-10 at various $l_{2}$ radii. On Imagenet, however, the latter gives better accuracy.
@@
Thus, the challenge of training a base classifier that is accurate and robust stands resolved [34] along with some noteworthy contributions [13] that ultimately prove that adversarial training is not a must for robust training and larger average certified radii can be achieved in less training time than state-of-the-art adversarial training algorithms.
III-B Improving Performance of Pre-Trained Classifiers
@@
Improving the performance of pre-trained classifiers by turning them from non-robust into provably robust ones is desired to maintain their confidentiality, integrity, or availability. For example, it might be beneficial to public vision API providers and users. In their work towards achieving this, Salman, Hadi, et al. [18] apply randomized smoothing to pre-trained classifiers by attaching a denoiser before the pre-trained model. Intuitively, they remove the Gaussian noise added in the randomized smoothing, unlike traditional methods that focus on increasing the quality of images. They train the denoiser with a modified loss function, which, instead of using the usual MSE, uses cross-entropy loss to train the combined denoiser and pre-trained classifier architecture, keeping the pre-trained classifiers’ weights frozen. This process also increases the classification accuracy. The methodology can be extended for other $l_{p}$ cases by simply changing the noise used during training to a noise sampled from the distribution corresponding to the specific $l_{p}$ case. Denoised smoothing can be applied to both white and black box settings, making it the state-of-the-art approach to defending pre-trained classifiers. Sheikholeslami, Fatemeh, et al. [54] challenge the low natural accuracy of this state-of-the-art approach and contribute further to the methodology by Salman, Hadi, et al. [18] by augmenting the joint system with a “rejector” and exploiting adaptive sample rejection, i.e., intentionally abstain from providing a prediction. The reject class is used, through Monte Carlo (MC) sampling, to reject noisy samples whose prediction is inconsistent with the prediction of the clean sample. This approach to preventing misclassified examples in the non-rejected classes has substantially improved the natural accuracy and the certification radius for computationally inexpensive denoisers. They achieve considerably better accuracy with their rejection methodology on CIFAR10 with both DnCNN-based denoising and MemNet-based denoising [110, 111]. Lee, Kyungmin, and Seyoon Oh. [46] also improve on previous works [18, 54] by introducing an improved score-based architecture for denoisers instead of classification loss, however, faced challenges concerning the novelty of their work and the possibility that the proposed method might exacerbate the weakness of randomized smoothing (i.e., slow prediction), especially in high-dimensions. Hence, while they produced higher certified accuracy of ResNet-110 on CIFAR-10 at various $l_{2}$ radii compared to previous works [1, 18], they failed to achieve similar improvements in certified accuracy of ResNet-50 on ImageNet, indicating that their method seems to be effective for low-resolution images only. While the work so far in this domain seems somewhat promising, there remains scope for future research to further improve the certified accuracy of high-dimensional datasets with increasing $\sigma$ ($\sigma$ as defined in section I, subsection C).
III-C Leveraging Local Information
The original randomized smoothing technique makes use of Monte-Carlo sampling to derive certified radii. This information is fully summarised for an input $x$ by $p_{A}(x)$, the probability of the top class, and does not require any other information beyond the standard deviation $\sigma$ of the Gaussian. Levine et al., [14] propose the use of additional gradient information, viz. $p_{A}^{\prime}(x)$, in a technique known as second-order smoothing (SoS) to provide more robust bounds. Their approach is based on the Lipschitz property of randomized smoothing classifiers [112], which globally bounds the Lipschitz constant of the gradient of a smoothed network. Thus, instead of using the Lipschitz bounds directly on the classifier, it utilizes the Lipschitz bounds on the gradient to obtain what the paper terms as second-order smoothing certificates, which can be used to compute a lower bound on the distance of a point to its closest adversarial example and improve robustness certificates. It also introduces Gaussian dipole smoothing, which is an optimisation based on SoS and Gaussian randomized smoothing to provide similar radius bounds. The results in this paper are significant as they show that all randomized smoothing-based classifiers suffer from this curvature constraint, restricting their application and prompting further research into certificates obtained using higher-order derivative information. This is what Mohapatra, Ko, et al. [24] do by extending the usage of gradient information to provide a general framework for evaluating more optimal certified radii using higher-order derivatives around the input. They do so by reformulating radius calculation as an optimisation problem. It also evaluates the certified radii obtained using Gaussian-smoothed classifiers, which utilize zeroth and first-order local information for general $l_{p}$ attacks. It also introduces a theoretical result for Gaussian-smoothed classifiers, which states that it is possible to achieve arbitrarily tight bounds using sufficient local information, thereby addressing the curse of dimensionality, which is a major challenge in RS and is discussed later on. Anderson et al. [20] utilize a locally optimal robust classifier based on Weierstrass transformations using decision boundary information. Their model not only improves on local information methods mentioned here but also addresses the issue of the robustness-accuracy trade-off for smoothing classifiers as discussed previously. Their study is limited to binary classifiers but remains unexplored for other cases, and this is a potential area for future works to explore.
@@ III-C Leveraging Local Information
Thus, one can see that there is an improvement in certified bounds using local information, such as the gradient and Hessian, and there is a possibility to explore and apply this approach orthogonally to other optimisations for randomized smoothing. A drawback of this approach is that the computational overheads associated with calculating local information add heavily to the already expensive smoothing convolution despite the gains being limited, as they also are tightly bound.
III-D Ensemble-based methods
@@
Horváth, Müller, et al. [57] theoretically motivate why ensembles are a particularly suitable choice as base models for RS and then empirically verify this choice, obtaining state-of-the-art results in multiple settings. The insight around which their work revolves is that reduced variance of ensembles over the perturbations introduced in RS leads to significantly more consistent classifiers. Additionally, they introduce a data-dependent adaptive sampling scheme for RS that enables a significant decrease in sample complexity of RS for predetermined radii, thus reducing its computational overhead. The modeling approach is mathematically related to the work by Tumer & Ghosh [91], which focuses on analyzing data and model choices to reduce classifier correlation for all ensemble methods, not just randomized smoothing.
RS requires augmenting data with large amounts of noise, which leads to a drop in accuracy. To overcome this, there is a training-free, modified smoothing approach, Smooth-Reduce, that leverages patching of images and aggregation to improve classifier certificates [27]. This approach is different from a simple ensemble technique, which is computationally expensive, and rather involves patch sampling. For this, the authors study two aggregation schemes and show that both approaches provide better certificates in certified accuracy compared to concurrent approaches. They also provide theoretical guarantees for such certificates and empirically show significant improvements over other methods that require expensive retraining and even extend this approach to videos.
@@
The Smoothed WEighted ENsembling (SWEEN) scheme improves the robustness of RS classifiers using ensemble techniques that generally can help achieve optimal certified robustness [45]. The authors make use of a $\gamma$-robustness index as a measure of optimality. Theoretical analysis further proves that the optimal SWEEN model can be obtained from training under mild assumptions. Additionally, SWEEN does not limit how individual candidate classifiers are trained. The model adopts a data-dependent weighted average of neural networks to serve as the base model for smoothing. To find the optimal SWEEN model, one must minimize the surrogate loss of $g_{ens}$ over the training set and obtain appropriate weights. The major drawback of ensembling is the high execution cost during inference, for which the paper also develops an adaptive prediction algorithm to reduce the prediction and certification cost of the models.
@@
From here, we see that ensemble-based methods for randomized smoothing have empirical and theoretical guarantees. Further, despite the prohibitive costs of ensemble training, adaptive algorithms for their training have been shown to reduce the computational overhead. Overall, ensemble methods provide a satisfactory approach to improve robustness guarantees with reduced costs as compared to other similarly expensive approaches. As with other approaches, we encourage the orthogonal application of these approaches and also a study on how to reduce ensemble training overheads further, possibly by advantageously using the noise augmentation already employed for randomized smoothing.
III-E Input Specific Information
@@
As discussed previously, RS uses a data-independent hyperparameter, $\sigma$, for the Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ to be convoluted with the base classifier. Even though this forms a basis of the robustness-accuracy tradeoff, it is also important to note that the issue of an optimal choice of $\sigma$ is important to improve certified robustness using TS. To deal with this, Alfarra et al. [15] propose the use of additional information, as is suggested in the curse of dimensionality section, which can be extracted from the data itself, in a scheme known as Data Dependent Randomized Smoothing (DDRS). DDRS aims to improve the choice of the hyperparameter $\sigma$ to achieve a larger certified radius, a parameter traditionally fixed in the original Randomized Smoothing (RS) method. Specifically, DDRS addresses this by solving an optimisation problem, $\sigma_{x}^{*}=\arg\max_{x}\;R(x,\mu)$.
This optimisation is performed using gradient ascent, and it defines the smoothed classifier G by introducing sampling noise $\mathcal{N}(0,{\sigma_{x}^{*}}^{2}I)$ for each input x. This approach has empirically demonstrated improved certified isotropic regions for individual data points. [44] also develop a related sample-wise choice of $\sigma$ which they term as Insta-RS along with adaptive training optimisations for the same. A major difference is that Insta-RS uses a multi-search, gradient-free paradigm, while DDRS assumes concavity and opts for gradient ascent.
@@
The concept of anisotropic smoothing aims to improve further on the results given by DDRS, which only certifies isotropic boundaries ([17], [115]). Anisotropic smoothing aims to certify, as the name suggests, different certified radii along each axis, essentially acting as a hyperelliptic curve instead of a hypersphere when considering the $l_{2}$ norm. Where normal randomized smoothing uses a constant variance noise of the form $\mathcal{N}\left(0,\sigma^{2}I\right)$ and DDRS uses a data-dependent variance of the form $\mathcal{N}\left(0,\sigma_{x}^{2}I\right)$ to certify robust bounds, anisotropic smoothing uses $\mathcal{N}\left(\mu,\Sigma\right)$ where $\Sigma$ is a diagonal matrix of variance generated using data-dependent techniques. Both papers follow a similar approach, differing only in the fact that in [115], $\mu=0$ is fixed, and they term their algorithm as ANCER. Their methodology is such that the certified bounds along each direction exceed not only the bounds presented in Cohen et al. [1], but also encapsulate the bounds given by isotropic data-dependent smoothing as in [15].
@@
In an unrelated but still data-dependent approach, Wang et al. [60] propose a smoothing algorithm that optimises the noise level choice (measured as the variance of the isotropic Gaussian noise) by dividing the input space into multiple regions and optimizing the noise level for each region, pre-training and fine-tuning the model using testing data. This approach allows one to create a sample-wise robust classifier that addresses the accuracy-robustness trade-off and improves the certified radii. Their approach maintains robustness in each region of the input space as per Cohen’s results without affecting other regions.
@@
Súkeník et al. [22] offers a concise overview of issues associated with a fixed global $\sigma$ value including (I) Using lower confidence bounds for class probability estimation, leading to smaller certified radii, (II) Balancing the trade-off between robustness and accuracy, as previously discussed and (III) The effect of Randomized Smoothing (RS) on the classifier’s decision boundary, where bounded or convex regions contract with increasing $\sigma$ while unbounded and anti-convex regions expand. The paper then tackles these issues by introducing Input-Specific Randomized Smoothing (ISRS) which employs the Neyman-Pearson lemma in a general case. However, they also state the effect of the curse of dimensionality and propose a framework to evaluate the efficacy of data-dependent RS methods in high dimensionality. It is a fundamental work in that it identifies the shortcomings of previous works like DDRS, Insta-RS, ANCER, and [60] which do have significant empirical results, but weak theoretical foundations.
@@
Riemannian data-dependent randomized smoothing (RDDRS) [38] further optimises the approach in anisotropic smoothing by allowing it to be robust to rotations. It does so using Riemannian optimisation to obtain $\mathcal{C}$, a covariance matrix that is not restricted to a scalar matrix (as in normal randomized smoothing and DDRS) nor a diagonal matrix (as in anisotropic randomized smoothing). Here, the smoothened classifier is constructed using Gaussian noise of the form $\mathcal{N}\left(0,\mathcal{C}\right)$. RANCER [116] also employs a non-axis-aligned approach. These methods empirically improve on the certified bounds introduced by even anisotropic smoothing techniques. As can be seen, there is clearly a pattern on increasing the complexity of the covariance matrix for the Gaussian noise to derive better results. As motivated by [22], however, we also note the weak theoretical justifications and encourage future work to analyse strongly the theoretical foundations to potentially justify and back the empirical results of approaches like DDRS, Insta-RS, ANCER, RDDRS, and RANCER.
Geometrically Informed Certified Robustness [5] is an approach which exploits the underlying geometric properties of a network using information around a point, such as the certified radii of nearby points, and exploits the transitivity of certifications. It is an optimisation that can be combined orthogonally with previous approaches, which only certify the region around a point without using any local information. We strongly suggest future empirical work to explore GICR in conjuction with other optimisations.
@@
Randomized smoothing has primarily been applied to single output cases. [61] propose a localised smoothing algorithm for multi-output settings instead of using independent RS schemes. Their approach combines anisotropic smoothing and also segments each input into various parts, where different anisotropic noise distributions are used for smoothing. In practice, they utilise soft locality, i.e. when the entire input influences the output but specific subsets have higher importance, to optimise the certified bounds and aggregate the result for the overall classifier. Their approach is a novel approach which extends and optimises RS for the multi-output scenario, and can be explored further, potentially with non-axis aligned distributions as well, similar to how RANCER extends ANCER.
@@
One of the issues mentioned in [22] is that of fairness where class boundary disparities affect classwise prediction accuracy and robustness. This also relates heavily to the discussion in robustness-accuracy tradeoff, where a paper highlighted the need to limit data geometry-dependent augmentation schemes for similar reasons [32]. Addressing this, Kumar et al. [21] use the average prediction score of a class and the margin by which the average prediction score of one class exceeds that of another. Following this, using a modified version of the Neyman-Pearson lemma, which forms the basis of RS and IDRS, a procedure is designed to compute the certified radius where the confidence is guaranteed to stay above a certain threshold, thus achieving a significantly better-certified radius, as is empirically shown. They raise a valid point regarding certifying confidence (soft classifiers) instead of labels (hard classifiers), as has been the primary focus of most RS based techniques so far. We urge further exploration and extension of this notion of understanding classifier confidence and base classifier properties and leveraging them for RS.
@@
Addressing the high cost challenge in RS, Quasi-Convexity-based Randomized Smoothing (QCRS) [59] presents an optimisation which exploits the quasiconcavity condition for the radius-variance curve in randomized smoothing, which is a weaker form of concavity. While DDR assumed concavity for the certified radii, empirically, it is seen that this is not the case despite it showing significant certified radii improvements. On the other hand, quasiconcavity is shown to hold for most data points. Based on this, QCRS is presented as a computationally efficient approach with less overheads compared to other input-specific inference-based optimisation methods. Further, it is an entirely orthogonal approach to many other optimisation techniques in randomized smoothing, and as with other orthogonal approaches, we encourage empirical testing with other methods.

### v2

@@ A.4 Randomized Smoothing for Training
The idea of incorporating randomized smoothing objectives in adversarial training ([49]) roots from the intuition that doing so will result in boosted performance and improved robustness. Further, addressing the challenge of training a good base classifier that is accurate and robust when smoothed led to significantly better results ([14]). [49] consider a smoothed classifier which is a generalization of equation (LABEL:eq2), to soft classifiers, namely, functions $f:\mathbb{R}^{d}\rightarrow P(y)$ where $P(y)$ is the set of probability distributions over $y$. They then train a hard classifier which essentially takes the argmax of the soft classifier, using adversarial examples generated by describing an adversarial attack against the smoothed soft classifier ([24] and [37]). This results in a boosted empirical robustness and substantially improved certifiable robustness using the certification method of [8]. [14] improved upon this by deriving a new regularized risk in which the regularization can ’adaptively encourage’ the training of the base classifier. It is computationally efficient and can be implemented in parallel with other empirical defense methods, under standard (non-adversarial) and adversarial training schemes. They also designed a new certification algorithm, T-CERTIFY, that can leverage the regularization effect to provide tighter robustness and lower bounds that hold high probability. [63] challenged these attack-dependent and time costly adversarial training techniques ([8] and [49]) by introducing an algorithm called MACER, which learns robust models without using adversarial training. It outperforms [14] by a significant margin in terms of Certified top-1 accuracy on CIFAR-10 at various $l_{2}$ radii. On Imagenet, however, the latter gives better accuracy.
@@
Thus, the challenge of training a base classifier that is accurate and robust stands resolved ([14]) along with some noteworthy contributions ([63]) that prove that adversarial training is not necessary for robust training and larger average certified radii can be achieved in less training time than state-of-the-art adversarial training algorithms. We can link these results to scalability and conclude that these methods make RS somewhat scalable in terms of time.
A.5 Improving Performance of Pre-Trained Classifiers
@@
Improving the performance of pre-trained classifiers by turning them from non-robust into provably robust ones is desired to maintain their confidentiality, integrity, and availability. For example, it might be beneficial to public vision API providers and users. In their work towards achieving this, [50] apply randomized smoothing to pre-trained classifiers by attaching a denoiser before the pre-trained model. Their methodology increases the classification accuracy and can be extended for other $l_{p}$ cases by simply changing the noise used during training to a noise sampled from the distribution corresponding to the specific $l_{p}$ case. Denoised smoothing can be applied to both white and black box settings, making it the state-of-the-art approach to defending pre-trained classifiers. [52] challenge the low natural accuracy of this state-of-the-art approach and contribute further to the methodology by [50] by augmenting the joint system with a “rejector” and exploiting adaptive sample rejection, i.e., intentionally abstain from providing a prediction. Their approach has substantially improved the natural accuracy and the certification radius for computationally inexpensive denoisers. They achieve considerably better accuracy with their rejection methodology on CIFAR10 with both DnCNN-based denoising and MemNet-based denoising ([65] and [55]). [29] also improved on previous works ([50] and [52]) by introducing an improved score-based architecture for denoisers instead of classification loss, however, faced challenges concerning the novelty of their work and the possibility that the proposed method might exacerbate the weakness of randomized smoothing (i.e., slow prediction), especially in high-dimensions. Hence, while they produced higher certified accuracy of ResNet-110 on CIFAR-10 at various $l_{2}$ radii compared to previous works ([8] and [50]), they failed to achieve similar improvements in certified accuracy of ResNet-50 on ImageNet, indicating that their method seems to be effective for low-resolution images only. While the work so far in this domain seems somewhat promising, there remains scope for future research to further improve the certified accuracy of high-dimensional datasets with increasing $\sigma$ ($\sigma$ as defined in section I, subsection C).
A.6 Leveraging Local Information
The original randomized smoothing technique makes use of Monte-Carlo sampling to derive certified radii. This information is fully summarised for an input $x$ by $p_{A}(x)$, the probability of the top class, and does not require any other information beyond the standard deviation $\sigma$ of the Gaussian. [32] propose the use of additional gradient information, viz. $p_{A}^{\prime}(x)$, in a technique known as second-order smoothing (SoS) to provide more robust bounds. Their approach is based on the Lipschitz property of randomized smoothing classifiers ([31]), which globally bounds the Lipschitz constant of the gradient of a smoothed network, showing that all randomized smoothing-based classifiers suffer from this curvature constraint, restricting their application and highlighting scalability issues. It also introduces Gaussian dipole smoothing as an optimisation based on SoS and Gaussian randomized smoothing to provide similar radius bounds, addressing the high inference cost issue.
[40] extend the usage of gradient information to provide a general framework for evaluating more optimal certified radii using higher-order derivatives around the input. They introduce a theoretical result for Gaussian-smoothed classifiers, which states that it is possible to achieve arbitrarily tight bounds using sufficient local information, thereby addressing the curse of dimensionality, which is a major challenge in RS.
[2] utilize a locally optimal robust classifier based on Weierstrass transformations using decision boundary information. Their model not only improves on local information methods mentioned here but also addresses the issue of the robustness-accuracy trade-off for smoothing classifiers as discussed previously. Their study is limited to binary classifiers but remains unexplored for other cases, and this is a potential area for future works to explore.
@@ A.6 Leveraging Local Information
Thus, one can see that there is an improvement in certified bounds using local information, and there is a possibility to explore and apply this approach orthogonally to other optimisations for randomized smoothing. However, from the perspective of scalability, a drawback of this approach is that the computational overheads associated with calculating local information add heavily to the already expensive smoothing convolution despite the gains being limited and tightly bound (as shown in [32]).
A.7 Ensemble-based methods
@@
[18] empirically verify choosing ensembles as base models for RS and obtain state-of-the-art results in multiple settings. The insight around which their work revolves is that reduced variance of ensembles over the perturbations introduced in RS leads to significantly more consistent classifiers. Additionally, they introduce a data-dependent adaptive sampling scheme for RS that enables a significant decrease in sample complexity of RS for predetermined radii, thus reducing its computational overhead. The modeling approach is mathematically related to the work by [57]. RS requires augmenting data with large amounts of noise, which leads to a drop in accuracy. Smooth-reduce, by [20], a training-free, modified smoothing approach overcomes this by leveraging the patching of images and aggregation to improve classifier certificates. This approach is different from a simple ensemble technique, which is computationally expensive, and instead involves patch sampling. They provide theoretical guarantees for such certificates and empirically show significant improvements over other methods. They also extend this approach to videos.
@@
The Smoothed WEighted ENsembling (SWEEN) scheme by [14] improves the robustness of RS classifiers using ensemble techniques that generally can help achieve optimal certified robustness. Theoretical analysis further proves that the optimal SWEEN model can be obtained from training under mild assumptions. Additionally, SWEEN does not limit how individual candidate classifiers are trained. The model adopts a data-dependent weighted average of neural networks to serve as the base model for smoothing. To find the optimal SWEEN model, one must minimize the surrogate loss of $g_{ens}$ over the training set and obtain appropriate weights. The major drawback of ensembling is the high execution cost during inference, for which the paper develops an adaptive prediction algorithm to reduce the prediction and certification cost of the models.
@@
Thus, we see that ensemble-based methods for randomized smoothing have empirical and theoretical guarantees. Despite the prohibitive costs of ensemble training, adaptive algorithms have been shown to reduce the computational overhead. However, these algorithms are built for a particular problem and may or may not optimize all kinds of ensemble models. Overall, ensemble methods provide a satisfactory approach to improve robustness guarantees with reduced costs as compared to other similarly expensive approaches. As with other approaches, we encourage the orthogonal application of these approaches and also a study on how to reduce ensemble training overheads further, possibly by advantageously using the noise augmentation already employed for randomized smoothing.
A.8 Input Specific Information
@@
RS uses a data-independent hyperparameter, $\sigma$, for the Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ to be convoluted with the base classifier. Even though this forms a basis of the robustness-accuracy tradeoff, it is also important to note that the issue of an optimal choice of $\sigma$ is important to improve certified robustness using TS. To deal with this, [1] propose the use of additional information, as is suggested in the curse of dimensionality section, which can be extracted from the data itself, in a scheme known as Data Dependent Randomized Smoothing (DDRS). DDRS aims to improve the choice of the hyperparameter $\sigma$ to achieve a larger certified radius, a parameter traditionally fixed in the original Randomized Smoothing (RS) method. Specifically, this is done by solving an optimisation problem using gradient ascent. [7] also develop a related sample-wise choice of $\sigma$ which they term as Insta-RS along with adaptive training optimisations for the same.
@@
The concept of anisotropic smoothing aims to improve further on the results given by DDRS, which only certifies isotropic boundaries ([17] and [11]). Anisotropic smoothing aims to certify, as the name suggests, different certified radii along each axis, essentially acting as an axis-aligned hyperelliptic curve instead of a hypersphere when considering the $l_{2}$ norm. Anisotropic smoothing uses $\mathcal{N}\left(\mu,\Sigma\right)$ where $\Sigma$ is a diagonal matrix of variance generated using data-dependent techniques as the noise distribution. Their methodology is such that the certified bounds along each direction exceed and encapsulate the bounds given by isotropic data-dependent smoothing as in [1].
@@
In an unrelated but still data-dependent approach, [60] propose a smoothing algorithm that optimises the noise level choice (measured as the variance of the isotropic Gaussian noise) by dividing the input space into multiple regions and optimizing the noise level for each region, pre-training and fine-tuning the model using testing data. This approach allows one to create a sample-wise robust classifier that addresses the accuracy-robustness trade-off and improves the certified radii. Their approach maintains robustness in each region of the input space as per Cohen’s results without affecting other regions.
@@
[53] offers a concise overview of issues associated with a fixed global $\sigma$ value including (I) Using lower confidence bounds for class probability estimation, leading to smaller certified radii, (II) Balancing the trade-off between robustness and accuracy, as previously discussed and (III) The effect of Randomized Smoothing (RS) on the classifier’s decision boundary, where bounded or convex regions contract with increasing $\sigma$ while unbounded and anti-convex regions expand. The paper then tackles these issues by introducing Input-Specific Randomized Smoothing (ISRS) which employs the Neyman-Pearson lemma in a general case. However, they also state the effect of the curse of dimensionality and propose a framework to evaluate the efficacy of data-dependent RS methods in high dimensionality.
@@
Riemannian data-dependent randomized smoothing (RDDRS) and RANCER ([26] and [48]) further optimises the approach in anisotropic smoothing by allowing it to be robust to rotations. It does so using Riemannian optimisation to obtain $\mathcal{C}$, a full covariance matrix for the Gaussian noise. As can be seen, there is clearly a pattern on increasing the complexity of the covariance matrix for the Gaussian noise to derive better results. As motivated by [53], however, we also note the weak theoretical foundations to justify and back the empirical results of approaches like DDRS, Insta-RS, ANCER, RDDRS, and RANCER.
@@
Geometrically Informed Certified Robustness by [10] is an approach which exploits the underlying geometric properties of a network using information around a point, such as the certified radii of nearby points, and exploits the transitivity of certifications. [51] propose a localised smoothing algorithm for multi-output settings instead of using independent RS schemes. Their approach segments each input into various parts using different anisotropic noise distributions. In practice, they utilise soft locality, i.e. when the entire input influences the output but specific subsets have higher importance, to optimise the certified bounds and aggregate the result.
@@
One of the issues mentioned in [53] is that of fairness where class boundary disparities affect classwise prediction accuracy and robustness. This also relates heavily to the discussion in robustness-accuracy tradeoff, where a paper highlighted the need to limit data geometry-dependent augmentation schemes for similar reasons [41]. Addressing this, citeb23 use a modified version of the Neyman-Pearson lemma to compute the certified radius where the confidence is guaranteed to stay above a certain threshold, thus achieving a significantly better-certified radius. [3] discuss randomized smoothing with the consideration of generic smoothing distributions, which can be input-dependent and provide an approach to solve the problem of maximising certified radii using input-dependent distributions. However, the paper mainly focuses on the theoretical side with a simple yet unpractical experiment as a proof of concept.
@@
Addressing the high cost challenge in RS, Quasi-Convexity-based Randomized Smoothing (QCRS) introduced by [23] presents an optimisation which exploits the quasiconcavity condition for the radius-variance curve in randomized smoothing, which is a weaker form of concavity. Quasiconcavity is shown to hold for most data points. Based on this, QCRS is presented as a computationally efficient approach with less overheads compared to other input-specific inference-based optimisation methods, one of the few input-specific methods to address complexity.